Solve for complete integral

z=px+qy+c√(1+p²+q²)


1
Expert's answer
2021-01-28T05:13:53-0500

This PDE is of Clariut's type. That is, it's of the form "z=px+qy+f(p, q)".

To get the complete integral of this kind of PDE, we replace "p, q" with "a, b" respectively.

Hence the complete integral of the PDE above is


"z=ax+by+c\sqrt{(1+a^2+b^2)}"